A bus has no legs. There are 7 girls in a bus. Each girl has 7 backpacks. In each backpack there are 7 big cats. For every big cat there are 7 little cats.
Question: How many legs are there in the bus?

Answer by edjones(8007)

7 little cats have 28 legs
1 big cat + 7 little cats = 32 legs
1 backpack= 7*32=224 legs
7 backpacks=7*224=1568 legs
7 girls=7(1568+2)=10,990 legs (I think)
.
Ed